A ladder 8ft long resting against a house makes a 60 degree angle with the ground.How far up the house does it reach? Give answer in terms with the radical in the numberator
Answer by stanbon(75887)

A ladder 8ft long resting against a house makes a 60 degree angle with the ground.How far up the house does it reach? Give answer in terms with the radical in the numerator.
-----------
Draw the picture.
You have a right triangle with hypotenuse = 8 ft
and ground angle = 60 degrees.
---
Find the height of the triangle.
Solve sin(60) = height/8
---
height = 8*sin(60) = 8*(sqrt(3)/2) = 4*sqrt(3) = 6.9282 ft.
=============================
Cheers,
Stan H.
============